OpenStudy (anonymous):

Write the equation of the quadratic function with roots -9 and and -3 and a vertex at (-6, -1).

OpenStudy (anonymous):

The formula I'm using is: f(x) = a[x^2 - (sum of roots)x + (product of roots)]

OpenStudy (triciaal):

roots -9 and -3 factors (x--9) and (x--3) a(x+9)(x+3)=0

OpenStudy (triciaal):

using (-6, -1) x = -6 and y = f(x) = -1 -1 =a (-6+9)(-6+3) a = -1/ (3* -3) = 1/9

OpenStudy (triciaal):